# Iodine and Bromine Radical Reactions in Atmospheric Mercury Oxidation

### v1

@@ III.1.4 Reaction R3 with X = Br or I and Y = OCl, OBr or OI
There is a general increase in bond length between oxygen and the middle halogen (HgX-OY) of 8% (0.16 Å) and a change in the XOY angle from 1.5 to 2 degrees between HgBrOY and HgIOY. Furthermore, there is a general trend that with the increasing atomic number of the halogens for the XHgOY and HgXOY molecules, the bond length and angle increase. When only increasing the atomic number of the end halogen (XHgO{Y} or HgXO{Y}), the bond length at the end of the complexes increases (X-HgOY, Hg-XOY / XHgO-Y, HgXO-Y), and the XHg-O-Y and HgX-O-Y bond angle increases as well, while the XHg-OY and HgX-OY bond length decreases. When the atomic number of the {X}HgOY or Hg{X}OY halogen increases, the XHgO-Y or HgXO-Y bong length also increases, while the (X-HgOY, XHg-OY / Hg-XOY, HgX-OY) bond lengths decrease together with the XHg-O-Y or HgX-O-Y bond angle. Geometries and frequencies can be found in the supplementary material section 1.3.
The largest difference in bond length is between HgBrOI and HgBrOCl of 15% (0.30 Å) for the HgXO-Y bond and the largest difference for the HgX-O-Y bond angle is between HgIOI and HgBrOI (4.6 degrees). The HgXOOH van der Waals complex differs only noticeably in the mercury halogen bond length (11% or 0.35 Å). There is also a small change in the HgXO-O-H bond angles, with a change of 2.23 degrees.
@@
The reactions forming the van der Waals complexes (reaction R3), follow the same trend with less negative reaction Gibbs free energies from HgXOI to HgXOCl, see Table 4. However, because of the characteristics of the HgX-OX bond, the HgBrOX forming reactions have less negative reaction Gibbs free energies than their HgIOX counterparts.
| Reaction R3 | X=I | X=Br | ||
| (298 K) | (0 K) | (298 K) | (0 K) | |
| Intermediate - Reactant (van der Waals complex) | ||||
| XHg+OI HgXOI | -28.37 | -38.43 | -21.88 | -31.65 |
| XHg+OBr HgXOBr | -16.30 | -25.97 | -9.89 | -19.23 |
| XHg+OCl HgXOCl | -13.81 | -23.32 | -7.45 | -16.65 |
| XHg+OOH HgXOOH | -6.45 | -16.50 | 0.72 | -10.34 |
| XHg+ONO HgXNO2 | -2.06 | -13.18 | 0.66 | -10.94 |
| Product - Intermediate (van der Waals complex) | ||||
| HgXOI Hg+XOI | -3.93 | 3.42 | -4.92 | 2.22 |
| HgXOBr Hg+XOBr | -3.79 | 3.54 | -5.79 | 1.11 |
| HgXOCl Hg+XOCl | -3.66 | 3.59 | -4.64 | 2.35 |
| HgXOOH Hg+XOOH | —– | —– | -4.54 | 3.38 |
| HgXNO2 Hg+XNO2 | -4.87 | 2.81 | -7.15 | 0.59 |
| Product - Reactant (van der Waals complex) | ||||
| XHg+OI Hg+XOI | -32.30 | -35.01 | -26.80 | -29.43 |
| XHg+OBr Hg+XOBr | -20.09 | -22.43 | -15.68 | -18.12 |
| XHg+OCl Hg+XOCl | -17.47 | -19.73 | -12.09 | -14.30 |
| XHg+OOH Hg+XOOH | —– | —– | -3.82 | -6.96 |
| XHg+ONO Hg+XNO2 | -6.93 | -10.37 | -6.49 | -10.35 |
III.1.5 Reaction R3 with X = Br or I and Y = NO or OOH
@@
Unlike to the case of HgBrOOH and HgBrNOO, where the dynamics of the reaction is changed from a positive energy difference to a reaction with a negative energy differnence, the HgINOO and HgIOOH van der Waals complexes only have negative energy differences, in the case of the reaction enthalpy at 0 K (-16.50 kcal/mol) and the Gibbs free energy at 298.15 K (-6.45 kcal/mol) for HgIOOH, or in the case of the reaction enthalpy at 0 K (-13.18 kcal/mol) and the Gibbs free energy at 298.15 K (-2.06 kcal/mol) for HgINOO.
If the reaction R3 for HgX + NO was the only reaction considered, then Hg would be more likely to be reformed in the HgBr + NO reaction than in the HgI + NO reaction. The same would apply to the HgBr + OOH reaction R3. In the other examples of the reversed reaction R3 investigated, this would also hold true, although the formation of Hg would be significantly more irrelevant and unlikely because of the van der Walls intermediates being significantly more stable.
@@ III.2 Kinetics
In the calculation of the reaction rates we have made use of a combination of CASSCF and CASPT2 calculations instead of the CCSD(T) approach employed in the calculation of the reaction energies and Gibbs free energies. Before we will analyse the results of these calculations we will first estimate the accuracy of these calculations by comparing to the CCSD(T) reaction energies.
@@ III.2.1 Estimation of Precision
The parameters that the Ktools program uses to calculate the rate coefficients are the enthalpy at zero Kelvin, the harmonic frequencies, and the rotational constants at many points along the reaction path. Therefore, a comparison between CASPT2 reaction energies extracted from the geometric scans, Figure 2, and CCSD(T) reaction energies, calculated from reactant and product energies, is made in Table 5.
@@
| HgBr | 12e12o | -70.9959 | -71.7103 | 0.7144 |
| HgI | 10e10o | -61.5792 | -60.9500 | -0.6292 |
| BrHgOBr | 10e10o | -56.2394 | -54.8740 | -1.3654 |
| BrHgOI | 12e12o | -57.9194 | -66.0154 | 8.0960 |
| IHgOBr | 14e14o | -51.0338 | -50.1649 | -0.8689 |
| IHgOI | 10e10o | -62.8126 | -63.3458 | 0.5332 |
| BrHgNOO | 16e13o | -34.8426 | -34.2926 | -0.5500 |
The electronic reaction energies calculated by the CASPT2 method with the ma-def2-TZVP basis set differ only slightly from the CCSD(T) electronic reaction energy. The two largest differences between the CCSD(T) and CASPT2 reaction energies are 1.37 kcal/mol for BrHgOBr and 8.10 kcal/mol for BrHgOI. The former is only slightly more than the chemical accuracy of 1 kcal/mol. As a rule of thumb, a difference of the order of the chemical accuracy in the activation energy leads to a factor of 5 in the rate coefficient. Therefore, the kinetic result for the reactions forming BrHgOI might not be as accurate as for the other reactions.
Another source of error concerns the precision of the frequency calculation. Calculating frequencies using the ECP relativistic approximation in the CASPT2 or CASSCF calculations resulted in too many imaginary frequencies. Therefore the ZORA relativistic approximation was used instead at the price of increasing the number of basis functions and thus expanding the size of the calculation. Obviously we cannot compare the ECP and ZORA calculated CASPT2 or CASSCF frequencies. Therefore, in Table 6, a comparison is made between the reaction energy calculated at the ECP CASPT2/ma-def2-TZVP level of theory and the ZORA CASPT2 or CASSCF/ma-def2-TZVP level of theory. This is to give an idea of the uncertainty in these calculations. The largest derivation of the energy is 16% in the case of IHgOI.
@@
| CASPT2 | 12e12o | HgBr | -70.50253 | -70.99589 | 1% |
| CASSCF | 10e10o | HgI | -59.2838 | -61.5792 | 4% |
| CASPT2 | 10e10o | BrHgOBr | -52.7045 | -57.8602 | 9% |
| CASSCF | 14e14o | IHgOBr | -62.1461 | -51.0338 | 10% |
| CASPT2 | 10e10o | IHgOI | -52.5597 | -62.8126 | 16% |
| CASPT2 | 10e10o | BHgOI | -55.3210 | -57.9194 | 5% |
@@ III.3 Rate Coefficient Results
The reaction rates in this study are calculated using the Multiwell program suite and its subprogram Ktools. The energy of the bond-dissociating reactions has been calculated using two different approaches, taking into account the distinct behavior of monoradical and biradical reactions. For reaction R1 calculations can be carried out using single reference methods, in this case CCSD(T). However, the energy for bond-breaking reactions forming biradicals is not described accurately at the CCSD(T) level of theory (reaction R2). Therefore, it has been necessary to use CASPT2 to accurately describe the dissociation energy,[64] although CCSD(T) reaction energies calculated from the equilibrium geometries of the product and reactants have been relied upon for comparison, Table 5. More information about the specifics of the CAS calculations can be found in the next section, and the input files for these calculations can be found in the supplementary material section 3.1.
@@
This article focuses on the difference between the reaction rates of iodine and bromine species. The kinetics of the reactions generally follow the trend as expected from the reaction energies, as can be seen in Table 7. The reactions containing bromine are generally faster. When comparing the formation of HgBr and HgI in reaction R2, there is a difference making the reaction forming HgBr twice as fast as the reaction forming HgI. Furthermore, the reaction forming BrHgOBr is about 7.2 times faster than the one forming IHgOI, about 6.5 times faster than that forming IHgOBr and about 10 times faster than that forming BrHgOI.
| Reactions | R1 | R2 | ||||||
| Product | HgBr | HgI | HgBr | HgI | BrHgOBr | IHgOI | IHgOBr | BrHgOI |
| Rate in | 3.310 | 5.510 | 8.810 | 4.210 | 3.310 | 4.610 | 5.110 | 3.310 |
| cmmolecules | ||||||||
Reaction |
rate coefficients |
Reference |
Hg+I→HgI |
4.0010 |
B3LYP/RRKM |
Skov 2004 [4] |
theory |
Hg + I → HgI |
1.27 (±0.58) 10 |
Raolfie 2008 [65] |
exp. |
Hg+Br→HgBr |
3.7010 |
CCSD(T)/RRKM |
Goodsite 2012 [66] |
theory |
9.810 |
Quas-CT |
Shepler 2007 [67] |
theory |
6.4(±.2)10 |
Sumner 2011 [68] |
exp. |
HgBr+Br→HgBr |
2.5010 |
B3LYP/RRKM |
Skov 2004 [4] |
theory |
3.0(±0.1)10 |
MRCI/Quasi-CT |
Balabanov 2005 [69] |
theory |
7.0010 |
at 397 K |
Greg 1970 [70] |
exp. |
Hg+Br→HgBr |
2.810 |
MRCI/ |
Balabanov 2005 [69] |
theory |
microcanonial VTST |
910 |
Ariya 2002 [71] |
exp. |
Hg+BrO→HgBrO |
(3.0–6.4)10 |
Spicer 2002 [72] |
exp. |
(1–100) 10 |
Raofie 2004 [73] |
exp. |
HgBr+NO→BrHgONO |
8.1410 |
PBE0// |
Jiao 2017 [56] |
theory |
RCCSDT/QuasCT |
Hg+NO→HgO+NO |
<410 |
Sommar 1997 [74] |
exp. |
<710 |
Sumner 2005 [75] |
exp. |
Hg+Cl→HgCl |
5.4010 |
Donohoue 2005 [76] |
exp. |
2.8010 |
CCSD(T)/CVTST |
Khalizov 2003 [54] |
theory |
Hg+ClO → HgClO |
1.110 |
Byun 2010 [77] |
exp. |
3.610 |
Spicer 2002 [72] |
exp. |
Hg+Cl → HgCl |
2.6(±0.2)10 |
Ariya 2002 [71] |
exp. |
2.5(±0.9)10 |
Sumner 2005 [75] |
exp. |
Hg + HCl → HgCl + H |
1.5010 |
QCISD |
Wilcox 2003 [78] |
theory |
1.0010 |
Hall 1993 [79] |
exp. |
Hg+F→HgF |
1.8610 |
CCSD(T)/CVTST |
Khalizov 2003 [54] |
theory |
Hg+OH→HgOH |
9.0(±1.3)10 |
Pal 2004 [80] |
exp. |
8.7(±2.8)10 |
Sommar 1997 [74] |
exp. |
HgOH+O→HgO+OH |
3.210 |
B3LYP/RRKM |
Skov 2004 [4] |
theory |
9.0(±1.3)10 |
Pal 2004 [80] |
exp. |
Hg+O→HgO+O |
7.5(±0.9)10 |
Pal 2004 [81] |
exp. |
3.110 |
QCISD(T)// |
Xu 2008 [82] |
theory |
MP2/TST |
Hg + HO → HgO +HO |
4.110 |
Seigneur 1994 [83] |
exp. |
<8.510 |
293 K |
Tokos 1998 [84] |
exp. |
| HgBr + OOH → BrHgOOH |
PBE0// |
Jiao 2017 [56] |
theory |
3.5110 |
RCCSDT/QuasCT |
@@ IV.1 Comparison with literature rate coefficients
Where possible, the rate coefficients calculated in this study by Ktools are compared with the rate coefficients presented in the literature. The values of the rates of reaction R1: X + Hg → HgX are in good agreement with the values shown in Table 8. They only differ slightly, with a factor of 0.89 for the reaction containing bromine and a factor of 1.38 for the reaction containing iodine compared to calculations made by Goodsite[66] and Skov.[4]
The forward rate coefficients given here are calculated by Ktools as the reverse rate coefficient divided by the equilibrium constant. When the reverse rate coefficient for reaction R2: HgBr → Br + HgBr is compared with the one calculated by Balabanov et al.[69] (5.510 s), the reverse rate coefficient calculated in this study (3.510 s) differs only by a factor of 0.63. The equilibrium constant given in Balabanov’s article is 5.40732510 moleculecm, which differs by only a factor of 1.4 from the equilibrium constant calculated in this study 3.910 moleculecm. The forward rate coefficient given in Balabanov’s article is 3.010 cmmolecules. The difference between the forward rate coefficient calculated in this study (8.810 cmmolecules) and the forward rate coefficient from Balabanov’s article would then differ by a factor of 34. However, if a forward rate coefficient is calculated as the reverse rate coefficient divided by the equilibrium constant, both from Balbanov’s article [69], the result is 1.01710 cmmolecules, which differs from the one calculated by Ktools only by a factor of 1.16. The rate coefficients calculated here are all calculated using the same approach, and are therefore suitable for the comparisons presented here.
The most influential difference of the result in this article is the result of reaction R2: X + HgX HgX, both because the rate of the reaction is the fastest calculated in this study and due to the similarities of the rates for the reactions, including bromine and iodine. The rate of reaction R2, including bromine, is close to the rate of the HgBr+NO→BrHgONO reaction with a rate of 8.110 cmmolecules and the HgBr+OOH→BrHgOOH reaction with a rate of 3.510 cmmolecules see Table 9.
@@
| species | cmmoleculess | species | cmmoleculess | Rate | ||
| HgBr | 8.810 | devided by | HgI | 4.210 | = | 2.09 |
| BrHgOBr | 3.310 | devided by | IHgOI | 4.610 | = | 7.20 |
| BrHgOBr | 3.310 | devided by | IHgOBr | 5.110 | = | 6.53 |
| BrHgOBr | 3.310 | devided by | BrHgOI | 3.310 | = | 10.0 |
| IHgOBr | 5.110 | devided by | IHgOI | 4.610 | = | 1.10 |
| IHgOBr | 5.110 | devided by | BrHgOI | 3.310 | = | 1.53 |
| IHgOI | 4.610 | devided by | BrHgOI | 3.310 | = | 1.39 |
For reaction R2: (OX + HgX XHgOX) the rate of the reactions is in the range between 3.310 - 3.310 cmmolecules. The rate coefficient is about 7.2 times higher for reaction R2 including BrHg and OBr than the one including IHg and OI. The rate coefficients for reactions forming the products IHgOI, IHgOBr and BrHgOI do not largely differ. The rates for the reaction BrO + HgBr BrHgOBr have about 12.7 times smaller rate coefficient as IHg + I IHgI. Since the concentrations of OBr would generally be assumed larger than the concentration of I radicals, the BrHgOBr reaction R2 could be significant compared with the reaction forming IHgI. The concentration of the OI species must be about 10 times higher than that of the OBr radicals to compete and up to 18 thousand times higher than NO, OOH, and Br to compete with their reaction rates of the HgBr+NOBrHgONO, HgBr+OOHBrHgOOH, and Br + HgBr HgBr reactions.
To further the understanding of how iodine competes with bromine reactions in the atmosphere the reactions of HgI+NO→IHgONO and HgI+OOH→IHgOOH need further investigation. From reaction energies, the bromine reactions would be favored over the iodine reactions, but certainty cannot be established without further investigation of the kinetics.
From the literature, Table 8, there are known reactions, including halogens and mercury that have rate coefficients in a range faster than BrHg + OBr BrHgOBr (3.310cmmolecules). These are:
@@
HgOH + O→HgO(s) + OH
Because the overall rate of the reactions is dependent both on the rate coefficient and the atmospheric concentrations of the chemical species, a brief estimate of the mixing ratio in the Arctic and Antarctic is presented in the next section.

### v2

@@ III.1.4 Reaction R3 with X = Br or I and OY = OCl, OBr or OI
The reactions forming the van der Waals complexes (reaction R3), follow the same trend with less negative reaction Gibbs free energies from HgXOI to HgXOCl, see Table 4. However, because of the characteristics of the HgX-OY bond, the HgBrOY forming reactions have less negative reaction Gibbs free energies than their HgIOY counterparts.
| Reaction R3 | X=I | X=Br | |||
| (298 K) | (0 K) | (298 K) | (0 K) | ||
| Intermediate - Reactant (van der Waals complex) | |||||
| XHg+OI HgXOI | -28.37 | -38.43 | -21.88 | -31.65 | |
| XHg+OBr HgXOBr | -16.30 | -25.97 | -9.89 | -19.23 | |
| XHg+OCl HgXOCl | -13.81 | -23.32 | -7.45 | -16.65 | |
| XHg+OOH HgXOOH | -6.45 | -16.50 | 0.72 | -10.34 | |
| XHg+ONO HgXNO2 | -2.06 | -13.18 | 0.66 | -10.94 | |
| Product - Intermediate (van der Waals complex) | |||||
| HgXOI Hg+XOI | -3.93 | 3.42 | -4.92 | 2.22 | |
| HgXOBr Hg+XOBr | -3.79 | 3.54 | -5.79 | 1.11 | |
| HgXOCl Hg+XOCl | -3.66 | 3.59 | -4.64 | 2.35 | |
| HgXOOH Hg+XOOH | —– | —– | -4.54 | 3.38 | |
| HgXNO2 Hg+XNO2 | -4.87 | 2.81 | -7.15 | 0.59 | |
| Product - Reactant (van der Waals complex) | |||||
| XHg+OI Hg+XOI | -32.30 | -35.01 | -26.80 | -29.43 | |
| XHg+OBr Hg+XOBr | -20.09 | -22.43 | -15.68 | -18.12 | |
| XHg+OCl Hg+XOCl | -17.47 | -19.73 | -12.09 | -14.30 | |
| XHg+OOH Hg+XOOH | —– | —– | -3.82 | -6.96 | |
| XHg+ONO Hg+XNO2 | -6.93 | -10.37 | -6.49 | -10.35 | |
III.1.5 Reaction R3 with X = Br or I and OY = NO2 or OOH
@@
Unlike to the cases of HgBrOOH and HgBrNOO, where the dynamics of the reaction are changed from a positive energy difference to a reaction with a negative energy differnence, the HgINOO and HgIOOH van der Waals complexes only have negative energy differences, in the case of the reaction enthalpy at 0 K (-16.50 kcal/mol) and the Gibbs free energy at 298.15 K (-6.45 kcal/mol) for HgIOOH, or in the case of the reaction enthalpy at 0 K (-13.18 kcal/mol) and the Gibbs free energy at 298.15 K (-2.06 kcal/mol) for HgINOO.
If reaction R3 for HgX + NO2 was the only reaction considered, then Hg would be more likely to be reformed in the HgBr + NO2 reaction than in the HgI + NO2 reaction. The same would apply to the HgBr + OOH reaction R3. In the other examples of the reversed reaction R3 investigated, this would also hold true, although the formation of Hg would be significantly more irrelevant and unlikely because of the van der Walls intermediates being significantly more stable.
@@ III.2 Kinetics
In the calculation of the reaction rates we have made use of a combination of CASSCF and CASPT2 calculations instead of the CCSD(T) approach employed in the calculation of the reaction energies and Gibbs free energies. Before we will analyse the results of these calculations we will first estimate the accuracy of these calculations by comparing to the CCSD(T) reaction energies. The bond dissociation curves calculated with the CAS methods can be seen in Figure 2.
@@ III.2.1 Estimation of Precision
To assess the accuracy of our calculations used to Ktools, we compared the CASPT2 reaction energies obtained from geometric scans with the CCSD(T) reaction energies calculated from reactant and product energies, as shown in Table 5.
@@
| HgBr2 | 12e12o | -71.00 | -71.71 | 0.71 |
| HgI2 | 10e10o | -61.58 | -60.95 | -0.63 |
| BrHgOBr | 10e10o | -56.24 | -54.87 | -1.37 |
| BrHgOI | 12e12o | -57.92 | -66.02 | 8.10 |
| IHgOBr | 14e14o | -51.03 | -50.16 | -0.87 |
| IHgOI | 10e10o | -62.81 | -63.35 | 0.53 |
| BrHgNO2 | 16e13o | -34.84 | -34.29 | -0.55 |
The electronic reaction energies calculated by the CASPT2 method differ only slightly from the CCSD(T) energies. The largest discrepancies are 1.37 kcal/mol for BrHgOBr and 8.10 kcal/mol for BrHgOI. While the former is close to the chemical accuracy of 1 kcal/mol, the latter suggests that the kinetic results for reactions forming BrHgOI may be less reliable. As a rule of thumb, a difference of the order of the chemical accuracy in the activation energy leads to a factor of 5 in the rate coefficient.
Frequency calculations using the ECP relativistic approximation resulted in too many imaginary frequencies, so the ZORA approximation was employed despite increasing computational cost. Direct comparison between ECP and ZORA frequencies is not feasible, in this instance.
To estimate uncertainties due to different relativistic approximations, we compared reaction energies calculated using ECP CASPT2/ma-def2-TZVP and ZORA CASPT2 or CASSCF/ma-def2-TZVP levels of theory. The largest deviation observed was 16% for IHgOI, indicating potential uncertainties in our calculations (see Table 6).
@@
| CASPT2 | 12e12o | HgBr2 | -70.50 | -71.00 | 1% |
| CASSCF | 10e10o | HgI2 | -59.28 | -61.58 | 4% |
| CASPT2 | 10e10o | BrHgOBr | -52.70 | -57.86 | 9% |
| CASSCF | 14e14o | IHgOBr | -62.15 | -51.03 | 10% |
| CASPT2 | 10e10o | IHgOI | -52.56 | -62.81 | 16% |
| CASPT2 | 10e10o | BHgOI | -55.32 | -57.92 | 5% |
@@ III.3 Rate Coefficient Results
The reaction rates in this study are calculated using the Multiwell program suite and its subprogram Ktools. The energy of the bond-dissociating reactions has been calculated, taking into account the distinct behavior of biradical reactions. The energy for bond-breaking reactions forming biradicals is not described accurately at the CCSD(T) level of theory (reaction R2). Therefore, it has been necessary to use CASPT2 to accurately describe the dissociation energy, [85] although CCSD(T) reaction energies calculated from the equilibrium geometries of the product and reactants have been relied upon for comparison, Table 5. More information about the specifics of the CAS calculations can be found in the next section, and the input files for these calculations can be found in the supplementary material section 3.1.
@@
This article focuses on the difference between the reaction rates of iodine and bromine species. The kinetics of the reactions generally follow the trend as expected from the reaction energies, as can be seen in Table 7. The reactions containing bromine are generally faster. When comparing the formation of HgBr2 and HgI2 in reaction R2, there is a difference making the reaction forming HgBr2 twice as fast as the reaction forming HgI2. Furthermore, the reaction forming BrHgOBr is about 7.2 times faster than the one forming IHgOI, about 6.5 times faster than that forming IHgOBr and about 10 times faster than that forming BrHgOI.
| Reactions | R2 | ||||||
| Product | HgBr2 | HgI2 | BrHgOBr | IHgOI | IHgOBr | BrHgOI | |
| 273 K | 7.910-13 | 3.910-13 | 3.410-14 | 4.910-15 | 4.610-15 | 3.810-15 | |
| 298 K | 8.810-13 | 4.210-13 | 3.310-14 | 4.610-15 | 5.110-15 | 3.310-15 | |
| Reaction | rate coefficients | Reference | ||
| Hg+I→HgI | 4.0010-13 | B3LYP/RRKM | [26] | theory |
| Hg + I2 → HgI2 | 1.27 (±0.58) 10-19 | [55] | exp. | |
| IHg + I → HgI2 | 4.210-13 | CASPT2/VTST | this study | theory |
| Hg+Br→HgBr | 3.7010-13 | CCSD(T)/RRKM | [28] | theory |
| 9.810-13 | Quas-CT | [66] | theory | |
| 6.4(±.2)10-13 | [76] | exp. | ||
| HgBr+Br→HgBr2 | 2.5010-10 | B3LYP/RRKM | [26] | theory |
| 3.0(±0.1)10-11 | MRCI/Quasi-CT | [6] | theory | |
| 7.0010-17 | at 397 K | [29] | exp. | |
| 8.810-13 | CASPT2/VTST | this study | theory | |
| Hg+Br2→HgBr2 | 2.810-31 | MRCI/VTST | [6] | theory |
| 910-17 | Ariya 2002 [1] | exp. | ||
| Hg+BrO→HgBrO | (3.0–6.4)10-14 | [71] | exp. | |
| (1–100) 10-15 | [54] | exp. | ||
| HgBr+NO2→BrHgONO | 8.1410-11 | PBE0// | [37] | theory |
| RCCSDT/QuasCT | ||||
| Hg+NO3→HgO+NO2 | <410-15 | [70] | exp. | |
| <710-15 | [75] | exp. | ||
| Hg+Cl→HgCl | 5.4010-13 | [24] | exp. | |
| 2.8010-12 | CCSD(T)/CVTST | [39] | theory | |
| Hg+ClO → HgClO | 1.110-11 | [11] | exp. | |
| 3.610-17 | [71] | exp. | ||
| Hg+Cl2 → HgCl2 | 2.6(±0.2)1018 | [1] | exp. | |
| 2.5(±0.9)10-18 | [75] | exp. | ||
| Hg + HCl → HgCl2 + H | 1.5010-33 | QCISD | [82] | theory |
| 1.0010-19 | [30] | exp. | ||
| Hg+F→HgF | 1.8610-12 | CCSD(T)/CVTST | [39] | theory |
| Hg+OH→HgOH | 9.0(±1.3)10-14 | Pal 2004 [50] | exp. | |
| 8.7(±2.8)10-14 | [70] | exp. | ||
| HgOH+O2→HgO+OH | 9.0(±1.3)10-14 | [50] | exp. | |
| Hg+O2→HgO+O2 | 7.5(±0.9)10-19 | [49] | exp. | |
| 3.110-40 | QCISD(T)//MP2/TST | Xu 2008 [84] | theory | |
| Hg + H2O2 → HgO +H2O | 4.110-16 | [62] | exp. | |
| <8.510-19 | 293 K | [77] | exp. | |
| HgBr + OOH → BrHgOOH | PBE0// | [37] | theory | |
| 3.5110-11 | RCCSDT/QuasCT | |||
| BHg + OBr → BrHgOBr | 3.310-14 | CASPT2/VTST | this study | theory |
| IHg + OI → IHgOI | 4.610-15 | CASPT2/VTST | this study | theory |
| IHg + OBr → IHgOBr | 5.110-15 | CASPT2/VTST | this study | theory |
| BrHg + OI → BrHgOI | 3.310-15 | CASPT2/VTST | this study | theory |
@@ IV.1 Comparison with literature rate coefficients
Of the reactions it has only been possible to compare the rate coefficients calculated by Ktools for reaction R2: Br + HgBr → HgBr2 with the rate coefficients presented in the literature. The forward rate coefficients given here are calculated by Ktools as the reverse rate coefficient divided by the equilibrium constant.
When the reverse rate coefficient for reaction R2: HgBr2 → Br + HgBr is compared with the one calculated by Balabanov et al. [6] (5.510-39 s-1), the reverse rate coefficient calculated in this study (3.510-39 s-1) differs only by a factor of 0.63. The equilibrium constant given in Balabanov’s article is 5.40732510-27 moleculecm-3, which differs by only a factor of 1.4 from the equilibrium constant calculated in this study 3.9175410-27 moleculecm-3. The forward rate coefficient given in Balabanov’s article is 3.010-11 cm3molecule-1s-1. The difference between the forward rate coefficient calculated in this study (8.810-13 cm3molecule-1s-1) and the forward rate coefficient from Balabanov’s article would then differ by a factor of 34. However, if a forward rate coefficient is calculated as the reverse rate coefficient divided by the equilibrium constant, both from Balbanov’s article [6], the result is 1.01710-12 cm3molecule-1s-1, which differs from the one calculated by Ktools only by a factor of 1.16. The rate coefficients calculated here are all calculated using the same approach, and are therefore suitable for the comparisons presented here.
The most influential difference of the result in this article is the result of reaction R2: X + HgX HgX2, both because the rate of the reaction is the fastest calculated in this study and due to the similarities of the rates for the reactions, including bromine and iodine. The rate of reaction R2, including bromine, is close to the rate of the HgBr+NO2→BrHgONO reaction with a rate of 8.110-11 cm3molecule-1s-1 and the HgBr+OOH→BrHgOOH reaction with a rate of 3.510-11 cm3molecule-1s-1 see Table 9.
@@
| species | cm3molecules-1s-1 | species | cm3molecules-1s-1 | Rate | ||
| HgBr2 | 8.810-13 | devided by | HgI2 | 4.210-13 | = | 2.09 |
| BrHgOBr | 3.310-14 | devided by | IHgOI | 4.610-15 | = | 7.20 |
| BrHgOBr | 3.310-14 | devided by | IHgOBr | 5.110-15 | = | 6.53 |
| BrHgOBr | 3.310-14 | devided by | BrHgOI | 3.310-15 | = | 10.0 |
| IHgOBr | 5.110-15 | devided by | IHgOI | 4.610-15 | = | 1.10 |
| IHgOBr | 5.110-15 | devided by | BrHgOI | 3.310-15 | = | 1.53 |
| IHgOI | 4.610-15 | devided by | BrHgOI | 3.310-15 | = | 1.39 |
For reaction R2: (OY + HgX XHgOY) the rate of the reactions is in the range of 3.310-14 to 3.310-15 cm3molecule-1s-1. The rate coefficient is about 7.2 times higher for reaction R2 including BrHg and OBr than the one including IHg and OI. The rate coefficients for reactions forming the products IHgOI, IHgOBr and BrHgOI do not largely differ.
Since the concentrations of OBr are generally higher than those of I, the BrHgOBr reaction could be significant compared to IHgI formation. For OI to compete, its concentration must be about 10 times higher than OBr and up to 18 thousand times higher than those of NO2, OOH, and Br.
Further investigation is needed to understand how iodine competes with bromine in atmospheric reactions, particularly for HgI + NO2 IHgONO and HgI + OOH IHgOOH. While reaction energies favor bromine reactions, kinetic studies are necessary for certainty.
From the literature, Table 8, there are known reactions, including halogens and mercury that have rate coefficients in a range faster than BrHg + OBr BrHgOBr (3.310-14cm3molecule-1s-1). These are:
@@
HgOH + O2→HgO(s) + OH
Since overall reaction rates depend on both rate coefficients and atmospheric concentrations, we provide an estimate of mixing ratios in the local conditions of Arctic and Antarctic in the next section.